A company makes computer chips. Minor defects in the materials (silicon, etc.) are common, and the probability that any one chip is defective is 25%. Therefore, the company tests each chip to make sure it is good before it sends it to market. The test that the company uses is not perfect – even a bad chip will pass the test 10% of the time, and a good chip will fail 5% of the time. In the long run, what fraction of the computer chips that the company sells are defective?
Answer #1

Ans:

P(defective)=0.25

P(pass/defective)=0.10

P(fail/not defective)=0.05

P(pass)=P(pass/defective)*P(defective)+P(pass/not defective)*P(not defective)

=0.10*0.25+(1-0.05)*(1-0.25)

=0.7375

Company will sell only those chips,that pass the test,so we have to calculate:

P(defective/pass)=P(pass/defective)*P(defective)/P(pass)

=0.10*0.25/0.7375=0.0339